How can I solve this:
y(y-12)-15=y(y+5)-3(y+5)

Answer by stanbon(75887)

y(y-12)-15=y(y+5)-3(y+5)
---------------------
y^2 - 12y - 15 = y^2 + 5y - 3y - 15
--------------
-12y - 15 = 2y - 15
14y = 0
y = 0
===============
Cheers,
Stan H.
---------------